OpenStudy (anonymous):

find f'(-2) if f(x)=(x+2)(x+3)(x+4)^2 the otherqoustion i have is find the volume if the graph of y= √20/π e^x/2 is routated about the x-axis from x=ln(1/2) to x=0

OpenStudy (ranga):

f(x) = (x+2)(x+3)(x+4)^2 = (x^2 + 5x + 6) * (x + 4)^2 Use product rule to find f'(x). Then put x = -2.

OpenStudy (kc_kennylau):

This is a step-by-step proof of the formula for the volume by rotation: http://www.wyzant.com/resources/lessons/math/calculus/integration/finding_volume

OpenStudy (kc_kennylau):

Basically it's just \(\pi\int ydx\)

OpenStudy (anonymous):

thanks kc-kennylau